# LEDRO: LLM-Enhanced Design Space Reduction and Optimization for Analog Circuits

### v1

@@ I Introduction
Recent works explore utilizing LLMs for analog circuit design [8, 9, 10, 11, 12, 13], but most of them don’t focus on transistor level sizing optimization. Refs. [8, 9] utilize LLMs for analog layout design, [10] uses agentic LLMs to automate design process, and [11, 13] prioritize topology and correct functionality of analog circuits. Ref. [12] focuses on transistor sizing optimization by expecting LLMs to directly provide the optimized sizing value, which can be challenging to scale to larger search spaces. To overcome these shortcomings, we introduce a novel framework - LLM-Enhanced Design Space Reduction and Optimization (LEDRO) - that synergizes the mathematical reasoning from optimization techniques with the circuit knowledge of LLMs. Fundamentally, LEDRO enhances design space exploration by iteratively leveraging LLMs to choose a refined search region (instead of direct points) and optimization techniques to find high-performing circuits in the chosen region. Fig. 1 provides a high-level illustration of LEDRO. To ensure LEDRO’s generalizability across technology nodes and topologies, we perform calibration point synthesis using optimization techniques to provide reference examples to the LLM to generate refined regions without requiring any circuit-specific training. To further improve LLM outputs, we provide LLM with its knowledge-based self-reflection and simulation-based optimization feedback. In our work, we utilize LLaMa3-70B [14] as the LLM and TuRBO [15] as the optimization technique for LEDRO.
@@
The rest of this paper is organized as follows. Section II introduces the proposed LEDRO method with a case study example. Section III provides experiment results on a set of diverse circuits and Section IV concludes the paper.
@@ II Methods
Our proposed framework LEDRO utilizes a Large Language Model (LLM) in conjunction with an optimization module to more efficiently explore the search space for analog design. A circuit simulator simulates the created analog designs and estimates the performance. LEDRO continually explores and improves the design by running these components in iteration for a fixed set of rounds. In our work, we utilize LLaMa3-70B [14] as the LLM and TuRBO [15] as our optimizer. For LLaMa3-70B, we utilize greedy decoding with temperature of 0.8, and maximum generation length of 1000 tokens. In total, we run LEDRO for ten iterations. We provide a high-level block diagram of LEDRO in Fig 2 and a case study in II-D.
@@ II-A Calibration Point Synthesis for LLM
To provide these reference designs, we first sample 200 points by running the optimizer over the entire search space as allowed by the technology node. Next, we extract “good” points from these sampled points by filtering the design configurations for which the amplifier gain exceeds a threshold value. We set the threshold to a low 0 dB so that these “good” points are easy to find. In rare cases where no “good” points are found, the optimization can be run longer, or the threshold value can be lowered (although we didn’t face such an issue with our experiments). Finally, we rank these filtered “good” points on their Figure of Merit (described in the next paragraph) and add the top five designs to our calibrated initial prompt as the reference designs, as shown in Fig 1(A).
@@ II-D Case Study: End-to-end example of 7nm differential folded cascode amplifier
We present a case study for the design of a differential-ended folded cascode circuit (as shown in Fig. 5) in 7nm PTM-MG technology [22] using LEDRO. This circuit, chosen for its relevance in past literature [23, 24, 25], serves as a good benchmark for showcasing our model performance. We conduct calibration point synthesis with TuRBO and add them to the first round prompt for the LLM as shown in Fig 3. For subsequent rounds, we follow up with optimizer feedback and LLM’s self-reflection, as shown in Fig 4. We compare the optimization performance of two different runs of LEDRO utilizing two different LLMs with a pure TuRBO optimizer in terms of Gain, Unity Gain Bandwidth (UGBW), and FoM vs. optimizer steps in Fig 5. As observed, pure TuRBO over-optimizes UGBW and is unable to improve gain over the fixed 1200 iterations. Although TuRBO might be able to reach a similar optimal point, it would take many more optimization steps. On that hand, both runs of LEDRO achieve stronger gain, UGBW, and, in turn, high FoM in fewer optimizer steps. We conduct future comprehensive experiments only using LLaMa3-70B [14] since we validated the utility of GPT-3.5-turbo through this case study and since it is more expensive to run the GPT model.
@@ III-A Experimental Setup
Benchmarks: To demonstrate the broad generalizability of our framework, we conduct experiments using a diverse set of 22 amplifier topologies, including both PMOS and NMOS input configurations with single-ended and differential-ended output designs. To create this benchmark, we modify and repurpose the subcircuit annotation task in [26] to design the parametrized FinFET netlists. For richer benchmarking, we split the amplifiers into 10 low-complexity and 12 high-complexity Op-Amps based on the number of optimization parameters. To ensure the robustness of our study, we utilize predictive technology models for multi-gate transistors (PTM-MG) [22] across multiple technology nodes, specifically 7nm, 10nm, 14nm, and 16nm. We show the different Op-Amps along with their categorization and number of parameters in Fig 6.
@@
Baselines and Implementation: We utilize Cadence Spectre to run our simulations. Primarily, we compare our approach against the pure TuRBO optimizer [15] which is proposed for automated analog sizing in [3], and also used for subspace sampling in [19]. TuRBO is run for 1200 steps (BO-1200), matching our method’s maximum allowable evaluation budget (200 steps for calibration point synthesis and 100 optimizer steps $\times$ 10 LEDRO iterations). As a stronger baseline, we also compare with an over-optimized TuRBO run with an evaluation budget of 2000 steps (namely BO-2000). For a fair comparison, we ensure that LEDRO and the TuRBO baseline use the same starting points by continuing to train TuRBO using the initial 200 steps used for calibration point synthesis in LEDRO. We also include a Reinforcement Learning (RL) baseline, which was also run for 1200 and 2000 steps (namely RL-1200 and RL-2000). For this baseline, we modify the base model from [4], with the recent improvements of actor-critic networks of [16, 17]. We provide the actor with the region of operation knowledge and the critic with additional $v_{gs}$, $v_{ds}$, $g_{m}$, and $i_{ds}$ information for better optimization. Simulation-based approaches like genetic algorithms [7, 27] are sample-inefficient and poor in performance compared to TuRBO, as shown in [3, 4]. Also, the code for other LLM approaches like [10, 12] is not public and difficult to implement. Thus, we exclude such baselines for our comparisons.
@@ III-B Results
We show a detailed breakdown of the performance and efficiency for each amplifier topology and technology node in Fig 7 and Fig 8 for the low-complexity and high-complexity groups, respectively. We notice a clear trend wherein our model LEDRO not only achieves the best FoM performance but also reaches optimal points in lesser number of steps. Through Table 1, we present a summary of the average maximum FoM achieved across the low and high-complexity groups of amplifier designs across the different technology nodes. Our results show the 8-20% and 16-21% improvement in amplifier performance for low-complexity Op-Amps compared to the 2000-step TuRBO and 2000-step RL optimization, respectively, across technology nodes. These improvements are further amplified in the high-complexity Op-Amps with 41-55% improvement over TuRBO-2000 and 70-80% over RL-2000. Notably, this increased improvement from low to high-complexity is consistently observed across all technology nodes. Such consistency suggests that while low-complexity and low-dimensional circuits are easy to optimize even with more simplistic methods, designing higher-complexity circuits is still challenging for conventional optimization algorithms. On the other hand, our method LEDRO leverages LLMs to reduce and refine the high-complexity search space and efficiently use optimization on this lower-complexity refined space to provide strong improvements for the high-complexity circuits as well. Such results highlight the robustness, generalizability, and scalability of LEDRO.
We also compare the optimization speed in terms of number of steps required by LEDRO and the best baseline BO-2000 to achieve the best FoM achieved by the BO-2000 algorithm, as shown in Table I. Results show that LEDRO achieves an average speed-up of 1.54 $\sim$ 2.48$\times$ across low-complexity Op-Amps and 1.17 $\sim$ 2.38$\times$ for high-complexity Op-Amps across the technology nodes relative to the BO-2000 baseline. This demonstrates the superior optimization efficiency of LEDRO compared to traditional optimization techniques.
@@ IV Conclusion
In this paper, we introduce LEDRO, which leverages LLMs in synergy with optimization techniques to efficiently explore design space of Op-Amps. Our method advances analog circuit design automation by iteratively refining the design search space through the LLM’s prior knowledge and reasoning, thus facilitating strong optimization performance and better designs without human expertise. Experiments across a wide range of topologies and technology nodes demonstrate LEDRO’s generalizability with 13-48% FoM improvements and efficiency with 1.7-2.15$\times$ speed-ups compared to conventional optimization techniques. Furthermore, LEDRO is significantly stronger in designing high-complexity and high-dimensional circuits, which provides promise for more practical use cases. In conclusion, the generalizability and adaptability with other optimization techniques marks an advancement in the field, offering a solution for complex analog circuit design challenges.
@@ Acknowledgment
The work was supported by the MIT-IBM Watson AI Lab, and we thank them for their support.

### v2

@@ I Introduction
Recent works explore utilizing LLMs for analog circuit design [8, 9, 10, 11, 12, 13], but most of them do not focus on transistor level sizing optimization. Refs. [8, 9] utilize LLMs for analog layout design, [10] uses agentic LLMs to automate design process, and [11, 13] prioritize topology and correct functionality of analog circuits. Ref. [12] focuses on transistor sizing optimization by expecting LLMs to directly provide the optimized sizing value, which can be challenging to scale to larger search spaces. To overcome these shortcomings, we introduce a novel framework - LLM-Enhanced Design Space Reduction and Optimization (LEDRO) - that synergizes the mathematical reasoning from optimization techniques with the circuit knowledge of LLMs. Fundamentally, LEDRO enhances design space exploration by iteratively leveraging LLMs to choose a refined search region (instead of direct points) and optimization techniques to find high-performing circuits in the chosen region. Fig. 1 provides a high-level illustration of LEDRO. To ensure LEDRO’s generalizability across technology nodes and topologies, we perform calibration point synthesis using optimization techniques to provide reference examples to the LLM to generate refined regions without requiring any circuit-specific training. To further improve LLM outputs, we provide LLM with its knowledge-based self-reflection and simulation-based optimization feedback. In our work, we utilize LLaMa3-70B [14] as the LLM and TuRBO [15] as the optimization technique for LEDRO.
@@
The rest of this paper is organized as follows. Section II introduces the proposed LEDRO method with a case study example. Section III provides experiment results on a set of diverse circuits and Section IV concludes the paper. The code is available at https://github.com/dimplekochar/LEDRO.
@@ II Methods
Our proposed framework LEDRO utilizes a Large Language Model (LLM) in conjunction with an optimization module to more efficiently explore the search space for analog design. A circuit simulator simulates the created analog designs and estimates the performance. LEDRO continually explores and improves the design by running these components in iteration for a fixed set of rounds. In our work, we utilize LLaMa3-70B [14] as the LLM and TuRBO [15] as our optimizer. For LLaMa3-70B, we utilize greedy decoding with temperature of 0.8, and maximum generation length of 1000 tokens. In total, we run LEDRO for ten iterations. We provide a high-level block diagram of LEDRO in Fig. 2 and a case study in II-D.
@@ II-A Calibration Point Synthesis for LLM
To provide these reference designs, we first sample 200 points by running the optimizer over the entire search space as allowed by the technology node. Next, we extract “good” points from these sampled points by filtering the design configurations for which the amplifier gain exceeds a threshold value. We set the threshold to a low 0 dB so that these “good” points are easy to find. In rare cases where no “good” points are found, the optimization can be run longer, or the threshold value can be lowered (although we did not face such an issue with our experiments). Finally, we rank these filtered “good” points on their Figure of Merit (described in the next paragraph) and add the top five designs to our calibrated initial prompt as the reference designs, as shown in Fig 1(A).
@@ II-D Case Study: End-to-end example of 7nm differential folded cascode amplifier
We present a case study for the design of a differential-ended folded cascode circuit (as shown in Fig. 5) in 7nm PTM-MG technology [22] using LEDRO. This circuit, chosen for its relevance in past literature [23, 24, 25], serves as a good benchmark for showcasing our model performance. We conduct calibration point synthesis with TuRBO and add them to the first round prompt for the LLM as shown in Fig 3. For subsequent rounds, we follow up with optimizer feedback and LLM’s self-reflection, as shown in Fig 4. We compare the optimization performance of two different runs of LEDRO utilizing two different LLMs with a pure TuRBO optimizer in terms of Gain, Unity Gain Bandwidth (UGBW), and FoM vs. optimizer steps in Fig 5. As observed, pure TuRBO over-optimizes UGBW and is unable to improve gain over the fixed 1200 iterations. Although TuRBO might be able to reach a similar optimal point, it would take many more optimization steps. Both runs of LEDRO achieve stronger gain, UGBW, and, in turn, high FoM in fewer optimizer steps. We conduct future comprehensive experiments using only LLaMa3-70B [14] since we validated the utility of GPT-3.5-turbo through this case study and it is more expensive to run the GPT model.
@@ III-A Experimental Setup
Benchmarks: To demonstrate the broad generalizability of our framework, we conduct experiments using a diverse set of 22 amplifier topologies, including both PMOS and NMOS input configurations with single-ended and differential-ended output designs, ranging from 7 to 19 transistors. To create this benchmark, we modify and repurpose the subcircuit annotation task in [26] to design the parametrized FinFET netlists. For richer benchmarking, we split the amplifiers into 10 low-complexity and 12 high-complexity Op-Amps based on the number of optimization parameters. To ensure robustness, we utilize predictive technology models for multi-gate transistors (PTM-MG) [22] across multiple technology nodes, specifically 7nm, 10nm, 14nm, and 16nm. We show the different Op-Amps with their categorization and number of parameters in Fig 6.
@@
Baselines and Implementation: We utilize Cadence Spectre to run our simulations. Primarily, we compare our approach against the pure TuRBO optimizer [15] which is proposed for automated analog sizing in [3], and also used for subspace sampling in [19]. TuRBO is run for 1200 steps (BO-1200), matching our method’s maximum allowable evaluation budget (200 steps for calibration point synthesis and 100 optimizer steps $\times$ 10 LEDRO iterations). As a stronger baseline, we also compare with an over-optimized TuRBO run with an evaluation budget of 2000 steps (namely BO-2000). For a fair comparison, we ensure that LEDRO and the TuRBO baseline use the same starting points by continuing to train TuRBO using the initial 200 steps used for calibration point synthesis in LEDRO, and report the best of five runs. We also include a Reinforcement Learning (RL) baseline, which was also run for 1200 and 2000 steps (namely RL-1200 and RL-2000). For this baseline, we modify the base model from [4], with the recent improvements of actor-critic networks of [16, 17]. We provide the actor with the region of operation knowledge and the critic with additional $v_{gs}$, $v_{ds}$, $g_{m}$, and $i_{ds}$ information for better optimization. Simulation-based approaches like genetic algorithms [7, 27] are sample-inefficient and poor in performance compared to TuRBO as shown in [3, 4]. AmpAgent’s[10] device sizing agent primarily uses TuRBO algorithm. Since its code is not public, initial-value estimates were not available. Our method will show clear improvements over TuRBO, benchmarking against AmpAgent’s approach to our best capability. Despite efforts to replicate ADO-LLM’s [12] process on our case study, it yielded suboptimal results. Due to unavailable implementation details, we do not include this baseline for our comparisons.
@@ III-B Results
We show a detailed breakdown of the performance and efficiency for each amplifier topology and technology node in Fig 7 and Fig 8 for the low-complexity and high-complexity groups, respectively. We notice a clear trend wherein our model LEDRO not only achieves the best FoM performance but also reaches optimal points in lesser number of steps. Through Table I, we present a summary of the average maximum FoM achieved across the low and high-complexity groups of amplifier designs across the different technology nodes. Our results show the 8-20% and 16-21% improvement in amplifier performance for low-complexity Op-Amps compared to the 2000-step TuRBO and 2000-step RL optimization, respectively, across technology nodes. These improvements are further amplified in the high-complexity Op-Amps with 41-55% improvement over TuRBO-2000 and 70-80% over RL-2000. Notably, this increased improvement from low to high-complexity is consistently observed across all technology nodes. Such consistency suggests that while low-complexity and low-dimensional circuits are easy to optimize even with more simplistic methods, designing higher-complexity circuits is still challenging for conventional optimization algorithms. On the other hand, our method LEDRO leverages LLMs to reduce and refine the high-complexity search space and efficiently use optimization on this lower-complexity refined space to provide strong improvements for the high-complexity circuits as well. Such results highlight the robustness, generalizability, and scalability of LEDRO.
We also compare the optimization speed in terms of number of steps required by LEDRO and the best baseline BO-2000 to achieve the best FoM achieved by the BO-2000 algorithm, as shown in Table I. Results show that LEDRO achieves an average speed-up of 1.54 $\sim$ 2.48$\times$ across low-complexity Op-Amps and 1.17 $\sim$ 2.38$\times$ for high-complexity Op-Amps across the technology nodes relative to the BO-2000 baseline. To compare computational latencies of BO-2000 and LEDRO, we compare the average time to achieve the best FoM achieved by the BO-2000 algorithm, while also accounting for the LLM prompt-response time. Relative to BO-2000, LEDRO achieves this optimum in: 0.49 $\sim$ 0.72$\times$ time for low-complexity Op-Amps 0.47 $\sim$ 0.98$\times$ time for high-complexity Op-Amps, across technology nodes. This demonstrates the superior optimization efficiency of LEDRO compared to traditional optimization techniques.
@@ IV Conclusion
In this paper, we introduce LEDRO, which leverages LLMs in synergy with optimization techniques to efficiently explore design space of Op-Amps. Our method advances analog circuit design automation by iteratively refining the design search space through the LLM’s prior knowledge and reasoning, thus facilitating strong optimization performance and better designs without human expertise. Experiments across a wide range of topologies and technology nodes demonstrate LEDRO’s generalizability with 13-48% FoM improvements and efficiency with 1.7 $\sim$ 2.15$\times$ speed-ups compared to conventional optimization techniques. Furthermore, LEDRO is significantly stronger in designing high-complexity and high-dimensional circuits, which provides promise for more practical use cases. In conclusion, the generalizability and adaptability with other optimization techniques marks an advancement in the field, offering a solution for complex analog circuit design challenges.
@@ Acknowledgment
We thank MIT Microsystems Technology Laboratory (MTL) for compute and simulation environment access, and IBM BAM for LLM API access.